explain please

2. Which one of the following DE is exact? a. (x+y)dx+(xy+1) dy=0 b (e + y)<x+ſe+x)dy = 0 c.(ye* +1) dx +(e + xy) dy = 0 d.
Answer #1

- Date: is is 1 of the form exact it Given Differential ton - M da + N dy o om - ƏN oy ax. We check I a) (x+y) dat(xy+1) dy=0sy aa Hence am # ON Hence dy is not Exact. ex (e?y t) dx + Cexy + 2) dy=0 am = eny (00 Bay ON = exy (4) ax am tan ay ax Hence4) Linear DE xtyl = 23 +9y2 ak dy = x3 + xy2 dy = dx x3 + x y z then k=3 given Linear DE is homogeneous dy = 23+242 212 daSutaible form of general solution of o is y = yg the у сеСо го 4 Аа? В с